# RePo: Language Models with Context Re-Positioning

### v1

@@ Abstract
In-context learning is fundamental to modern Large Language Models (LLMs); however, prevailing architectures impose a rigid and fixed contextual structure by assigning linear or constant positional indices. Drawing on Cognitive Load Theory (CLT), we argue that this uninformative structure increases extraneous cognitive load, consuming finite working memory capacity that should be allocated to deep reasoning and attention allocation. To address this, we propose RePo, a novel mechanism that reduces extraneous load via context re-positioning. Unlike standard approaches, RePo utilizes a differentiable module, $f_{\phi}$, to assign token positions that capture contextual dependencies, rather than replying on pre-defined integer range. By continually pre-training on the OLMo-2 1B backbone, we demonstrate that RePo significantly enhances performance on tasks involving noisy contexts, structured data, and longer context length, while maintaining competitive performance on general short-context tasks. Detailed analysis reveals that RePo successfully allocate higher attention to distant but relevant information, assign positions in dense and non-linear space, and capture the intrinsic structure of the input context. Our code is available at https://github.com/SakanaAI/repo.
@@ 1 Introduction
Figure 1: Overall performance on four evaluation dimensions.
@@
The emergence of large language models (LLMs) (brown2020language) has enabled a wide range of applications, including few-shot learning (brown2020language), retrieval-augmented generation (lewis2020retrieval; yao2023react), and agentic systems (schick2023toolformer; park2023generative). At the core of these applications is in-context learning (shen2024position), a form analogous to human working memory, where information within a limited context window is temporarily stored and processed to solve a task. Consequently, exploring how to effectively utilize context information has become a fundamental research line in the LLM era (wei2022chain; weston2023system; chen2023extending).
@@
Recent studies show that an LLM’s ability to leverage contextual information is strongly influenced by its position encoding scheme (vaswani2017attention; press2021train; su2024roformer). Most LLMs impose a fixed contextual structure by assigning tokens consecutive integer indices from 0 to L1L-1 (vaswani2017attention) or a constant index $a$ for all tokens (kazemnejad2023impact). These indices are then integrated into a model through position encoding functions, enforcing a rigid organization of context.
@@
Although fixed position assignments have become the de facto standard, they deviate from how human working memory processes information. According to Cognitive Load Theory (CLT), the capacity of working memory during problem solving can be consumed by costs arising from how information is organized and presented, referred to as extraneous load (sweller1994cognitive; paas2003cognitive). CLT studies suggest that humans can actively reduce this extraneous load by reorganizing context, e.g., grouping related information into meaningful chunks (miller1956magical) or removing irrelevant details from instructions (sweller1994cognitive). Since working memory capacity is fixed, the capacity saved through such reorganization can be reallocated to deeper reasoning processes associated with the the germane load, thereby improving problem-solving performance (sweller1994cognitive).
@@
However, the critical ability to reorganize and restructure contextual information (vaswani2017attention; yang2025qwen3; dubey2024llama) is absent from the architectural design of modern LLMs. From the perspective of CLT (sweller1994cognitive; paas2003cognitive), rigid linear or constant position structures can be interpreted as introducing additional extraneous load, which in turn impairs attention allocation and deeper contextual reasoning (i.e., germane processing). As a consequence, tasks that require strong long-range or fine-grained contextual dependencies, e.g., needle-in-a-haystack (NIAH) problems (kamradt2023needle) or question answering under highly diluted contexts (hsiehruler), exhibit notable performance degradation, mirroring the effects predicted by CLT under high extraneous load. Moreover, from a probabilistic standpoint, these position assignment strategies, essentially uniform distributions over fixed integer ranges, are the least informative organizations of context and therefore limit representational efficiency.
@@
We propose an internal mechanism for LLMs to reduce extraneous load by re-organizing the positions of tokens. We formalize this process, termed context Re-Positioning (RePo), as learning non-constrained position values based on information relevance of tokens, instead of using the fixed linear positions in prior work. To this end, we introduce a differentiable module $f_{\phi}$, which assigns a position value in continuous space for each token based on its hidden state. The $f_{\phi}$ can be independently learned for each attention head of an LLM. Trained on general data, $f_{\phi}$ learns to re-position tokens free from conventional constraints like monotonicity or integer values. The continuity of modern position encoding functions, e.g., RoPE (su2024roformer) and ALiBi (press2021train) methods, is key to the end-to-end optimization of $f_{\phi}$, as it allows these assigned positions to be integrated in a differentiable manner.
@@
We find that LLMs using the RePo method achieve consistent performance gains on tasks involving noisy context, structured data, and longer context. In our experiments, we continually pre-trained LLMs with the RePo method and several baselines based on the OLMo-2 1B model. Within the training context length, our RePo method outperforms other baselines by at least 6.24 and 1.16 points on noisy context and structured data tasks, respectively. In addition, when extending the testing context length to 16K tokens using the YaRN (peng2024yarn) method, our RePo method outperforms other baselines by at least 13.25 EM points on the QA and Needle-in-the-haystack (NIAH) tasks and 5.48 points on LongBench (bai-etal-2024-longbench). Alongside these performance gains, RePo achieves comparable results on extensive general benchmarks (wang2024mmlu; clark2018think; zellers-etal-2019-hellaswag), which are typically short and require little reorganization.
@@
To better understand whether our method aligns with the CLT-based motivation and where the performance gains of the RePo method come from, we conducted a series of detailed analyses. First, to evaluate how different methods handle long-range dependencies, we compared the attention distributions across methods, particularly focusing on their treatment of distant but relevant information. In the NIAH task (Section 5.1), we observed that RePo allocates more attention to the most critical ``needle'' tokens compared to the baseline methods, while directing less attention to the nearest ``query'' tokens. This behavior breaks the typical locality bias (yang2025rope; press2021train), dynamically adjusting based on the context. Second, the positions assigned by RePo exist in a denser, more non-linear space, which is critical for enhancing its generalization to longer contexts, such as extending from 4K to 16K tokens. Finally, an interesting finding is that RePo learns position patterns that resemble a hybrid of previous position assignment strategies, such as assigning constant positions $a\pm\epsilon$ (kazemnejad2023impact) or enforcing a monotonic position sequence (vaswani2017attention; yang2025rope; press2021train), within a given context span (Section 5.2). In our case study (Appendix D), we also found that the positions assigned by RePo capture the intrinsic structure of the input context (e.g., segmentation of few-shot examples).
@@ 2 Background
The position information of tokens in a context is critical for the attention mechanism of LLMs. The position of a given token is generally mapped into embeddings (su2024roformer; vaswani2017attention; li2025seqpe) or biases (press2021train; raffel2020exploring) through a position encoding module before the attention function. In this section, we will introduce the notations for attention mechanism and position encoding functions.
@@
We use the rotary position encoding (RoPE) (su2024roformer) as a specific example to illustrate the position assignment and encoding procedure commonly employed in LLMs (walsh2; yang2025qwen3; qwen2025qwen25technicalreport). First, each token $x_{i}$ is assigned an integer position index $i$. This positional information is then incorporated into the model through an encoding function $g_{\theta}:\mathbb{R}\rightarrow\mathbb{R}^{d\times d}$, which is a differentiable function that generates a rotation matrix. The parameter $\theta$ represents a pre-defined frequency vector, generally frozen during training. In RoPE, the function $g_{\theta}$ is directly applied to the query $\bm{q}_{i}$ and key $\bm{k}_{j}$ as follows:
@@
In many related works, strict linear position assignment (i.e., 0 to L1L-1) similar to RoPE has become the standard approach, with various encoding functions being explored (vaswani2017attention; press2021train; yang2025rope; li2025seqpe). One notable exception is the NoPE method (kazemnejad2023impact; yang2025rope), which omits both position assignment and position encoding. However, we demonstrate that this approach is equivalent to applying position encoding at a constant position $a$ for all tokens. Further discussion about the inter-connection between constant (e.g., NoPE) and linear position assignment (e.g., RoPE) strategies are provided in Appendix A.
@@ 3.1 Overview
The current linear context organization, which assigns consecutive integer indices to tokens, overlooks the internal structure of tokens based on their relevance. For instance, this limitation leads to noticeable performance degradation in tasks involving long-distance dependencies (hsiehruler), a problem analogous to the high extraneous load issue in Cognitive Load Theory (CLT).
The main goal of this work is to reduce the unnecessary cognitive cost raised by the oversimplified organization of the context, i.e., extraneous cognitive load, thereby conserving the finite working memory capacity for beneficial germane load, such as the attention mechanism. To this end, we propose a context re-positioning (RePo) module $f_{\phi}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, which is a light-weight neural network that assigns more appropriate positions of tokens, taking into account their relevance within a given context. The assigned positions by $f_{\phi}$ are defined in a continuous, non-linear space, and can thus be optimized jointly with LLMs when equipped with differentiable position encoding methods (vaswani2017attention; su2024roformer; press2021train). Notably, our RePo module $f_{\phi}$ is prior to position encoding, where the latter aims to map assigned positions into embeddings or biases.
@@ 3.2 Context Re-positioning
kazemnejad2023impact shows that position information may be entangled in original hidden states, so the first component is designed to extract the position representation from hidden states of tokens explicitly. In our implementation, we use a light-weight SwiGLU (shazeer2020glu) sub-layer to achieve this goal:
@@
The subsequent component assigns a new position value $z_{i}$ for token $x_{i}$ in each attention head. There are a variety of modeling strategies for processing $\bm{r}_{i}$ with full (vaswani2017attention) or limited (lecun1998convolutional) access to $\bm{r}_{<i}$, but we find that a linear transformation achieves comparable performance with lower latency:
@@
where the position encoding function $g_{\theta}$ is the same as that used in RoPE in Eq. 3. It is worth noting that our RePo is not restricted to RoPE and can be easily extended to all the differentiable position encoding methods (vaswani2017attention; press2021train; li2025seqpe). In practice, we apply the position representation module (Eq. 4) for each layer independently, and position assignment module (Eq. 5) for each attention head independently.
Table 1: Performance on noisy context. We evaluate on subsets of RULER with noisy context, i.e., Needle-in-the-haystack (NIAH), Question Answering (QA), Aggregation (AGG), and variable tracking (VT), within the training context length (i.e., 4K tokens). This uses recall for NAIH, AGG, and VT tasks, and EM for the QA task as the metrics. We use bold and underline for best and second best results.
Model NIAH QA AGG VT AVG.
RoPE 82.56 57.00 37.98 1.00 44.64
NoPE 74.59 49.00 22.45 12.20 39.56
R2N1 85.00 59.50 31.10 0.00 43.90
N2R1 80.00 58.00 32.75 27.00 49.44
RePo 88.25 61.00 35.05 38.40 55.68
@@
Since many position encoding methods are differentiable (su2024roformer; press2021train; li2025seqpe), we can train an LLM with RePo-based attention (Eq. 7) using backpropagation when equipped with such encodings. To balance efficiency and effectiveness, we apply the RePo method starting from the $l$-th layer (e.g., $l=5$) while keeping standard position encoding for the lower layers. This design choice is motivated by previous findings that the lower layers of LLMs primarily capture surface-level features that depend more on local information (tenney-etal-2019-bert), such as part-of-speech tagging and syntax, and thus benefit less from reorganization. An ablation study for the hyper-parameter is in Appendix B.3.
@@ 4 Experiments
Table 2: Performance for structured data. We evaluate on NLGraph and HybridQA datasets for graph and table data, respectively. We use exact match (EM) as metric. Bold and underline denote the best and the second best results, respectively.
Graph Table AVG.
RoPE 27.43 24.43 25.93
NoPE 29.90 23.52 26.71
R2N1 27.11 25.11 26.11
N2R1 25.42 23.86 24.64
RePo 29.03 26.70 27.87
Figure 2: Long-context Evaluation on RULER. YaRN (peng2024yarn) is used for all RoPE layers to extend the context. We observe consistent results on the more realistic benchmark LongBench in Table 3.
@@ 4.1 Setup
We use OLMo-2 (walsh2) developed by Allen Institute for AI as the backbone model, whose performance is comparable to Qwen-2.5 (qwen2025qwen25technicalreport). Its data, model weights, and code are fully open-sourced. Notably, OLMo-2 is trained with the RoPE method (su2024roformer). For all methods in our experiments, we start from the checkpoint of the OLMo-2 1B model that has completed stage-1 pre-training, and we continually pre-train it on the 5050B-token stage-2 data [^3] https://github.com/allenai/OLMo. The training context length is 4096 tokens. We keep the training configuration and codebase identical to those released by walsh2.
@@
For our RePo method, we apply it starting from the 5th layer of the OLMo-2 1B model. In each layer that uses RePo, we share the parameters for position representation transformation in Eq. 4, while learning the position assignment for each attention head independently as in Eq. 5. The hidden size for the learned position representation is 256, i.e., $1/8$ of the model’s hidden size. As shown in §5.4, RePo is lightweight and introduces negligible overhead to the original LLM. We compare RePo with baselines:
  • RoPE: Uses RoPE (su2024roformer) for positional encoding, identical to pre-training.
@@
  • NoPE: Removes explicit positional encoding methods (kazemnejad2023impact; wang-etal-2024-length), i.e., RoPE is omitted during the continual pre-training.
@@
  • R2N1: A hybrid positional encoding method that interleaves RoPE and NoPE (yang2025rope; meta2025llama). For every three layers, the first two use RoPE while the last one uses NoPE.
  • N2R1: The opposite of R2N1, i.e., every three layers, the first two use NoPE and the last one uses RoPE.
@@
We train those models on 4 H100 GPUs for 50B tokens. We use the allenai/olmes [^4] https://github.com/allenai/olmes codebase for evaluation, which provides extensive test suites. We categorize our main evaluation tasks into three dimensions as follows:
@@
  • Noisy Context evaluates the model’s robustness to contexts containing large amounts of irrelevant information. Such ``noise’’ increases extraneous cognitive load (paas2003cognitive), which can negatively affect problem-solving performance. We use the RULER benchmark (hsiehruler), which purposefully constructs contexts with irrelevant content for evaluation.
@@
  • Structured Data evaluates performance on structured data such as tables and graphs. This setting highlights the importance of contextual organization, as linearizing such data into natural language often leads to significant structural information loss. We use NLGraph (wang2023can) and HybridQA (chen-etal-2020-hybridqa) to evaluate model performance on graph and table reasoning, respectively.
@@
  • Longer Context evaluates model performance on sequences longer than those seen during training (i.e., 4K tokens), as positional encoding has been shown to strongly affect long-context generalization (press2021train). We use subsets of RULER and LongBench (bai-etal-2024-longbench), which contain examples with 4K to 16K tokens for this evaluation.
@@
Notably, both the evaluations for noisy context and structured data are conducted within the training context length (i.e., 4K tokens). In contrast, the longer-context evaluation requires extended context lengths. To enable this, we apply the YaRN method to RoPE layers across all methods using the extrapolation hyperparameters in peng2024yarn. More details are in Appendix B.
@@
Table 3: Performance on LongBench. This uses F1 score for QA and fewshot tasks, while Rouge-L (lin-2004-rouge) for the summarization tasks. We only evaluate on data that contains less than 16K tokens. Since the model is trained with a max of 4096 tokens, the YaRN method (peng2024yarn) is used for all RoPE layers for context extrapolation. We use bold and underline for best and second best results, respectively.
@@
Method Multidoc QA Singledoc QA Summarization Fewshot AVG.
2WikiMultihopQA MuSiQue MultiFieldQA-en NarrativeQA GovReport QMSum TREC TriviaQA
RoPE 23.32 7.24 27.37 12.94 6.23 7.96 22.00 61.47 21.07
NoPE 9.11 0.33 13.64 1.80 5.15 0.68 9.50 18.12 7.29
R2N1 25.88 7.31 31.28 16.24 4.53 8.74 22.00 66.67 22.83
N2R1 16.24 0.40 21.88 1.26 8.74 5.31 21.50 25.18 12.56
RePo 30.86 13.45 33.12 15.24 16.80 12.53 31.50 73.02 28.31

### v2

@@ Abstract
In-context learning is fundamental to modern Large Language Models (LLMs); however, prevailing architectures impose a rigid and fixed contextual structure by assigning linear or constant positional indices. Drawing on Cognitive Load Theory (CLT), we argue that this uninformative structure increases extraneous cognitive load, consuming finite working memory capacity that should be allocated to deep reasoning and attention allocation. To address this, we propose RePo, a novel mechanism that reduces extraneous load via context re-positioning. Unlike standard approaches, RePo utilizes a differentiable module, $f_{\phi}$, to assign token positions that capture contextual dependencies, rather than replying on pre-defined order. By continually pre-training on the OLMo-2 1B & 7B models, we demonstrate that RePo consistently enhances performance on tasks involving noisy contexts, structured data, and longer context length, while maintaining competitive performance on general short-context tasks. Detailed analysis reveals that RePo successfully allocate higher attention to distant but relevant information, assign positions in dense and non-linear space, and capture the intrinsic structure of the input context. We will open-source the code and model weights. Our code is at https://github.com/SakanaAI/repo.
@@ 1 Introduction
The emergence of large language models (LLMs) (Brown et al., 2020) has enabled a wide range of applications, including few-shot learning (Brown et al., 2020), retrieval-augmented generation (Lewis et al., 2020; Yao et al., 2023), and agentic systems (Schick et al., 2023; Park et al., 2023). At the core of these applications is in-context learning (Shen et al., 2024), a form analogous to human working memory, where information within a limited context window is temporarily stored and processed to solve a task. Consequently, exploring how to effectively utilize context information has become a fundamental research line in the LLM era (Wei et al., 2022; Weston and Sukhbaatar, 2023; Chen et al., 2023).
@@
Recent studies show that an LLM’s ability to leverage contextual information is strongly influenced by its position encoding scheme (Vaswani et al., 2017; Press et al., 2021; Su et al., 2024). Most LLMs impose a fixed contextual structure by assigning tokens consecutive integer indices from 0 to L1L-1 (Vaswani et al., 2017) or a constant index $a$ for all tokens (Kazemnejad et al., 2023). These indices are then integrated into a model through position encoding functions, enforcing a rigid organization of context.
@@
Although fixed position assignments have become the de facto standard, they deviate from how human working memory processes information. According to Cognitive Load Theory (CLT), the capacity of working memory during problem solving can be consumed by costs arising from how information is organized and presented, referred to as extraneous load (Sweller, 1994; Paas et al., 2003). CLT studies suggest that humans can actively reduce this extraneous load by reorganizing context, e.g., grouping related information into meaningful chunks (Miller, 1956) or removing irrelevant details from instructions (Sweller, 1994). Since working memory capacity is fixed, the capacity saved through such reorganization can be reallocated to deeper reasoning processes associated with the germane load, thereby improving problem-solving performance (Sweller, 1994).
@@
However, the critical ability to reorganize and restructure contextual information (Vaswani et al., 2017; Yang et al., 2025a; Dubey et al., 2024) is absent from the architectural design of modern LLMs. From the perspective of CLT (Sweller, 1994; Paas et al., 2003), rigid linear or constant position structures can be interpreted as introducing additional extraneous load, which in turn impairs attention allocation and deeper contextual reasoning (i.e., germane processing). As a consequence, tasks that require strong long-range or fine-grained contextual dependencies, e.g., needle-in-a-haystack (NIAH) problems (Kamradt, 2023) or question answering under highly diluted contexts (Hsieh et al., ), exhibit notable performance degradation, mirroring the effects predicted by CLT under high extraneous load.
@@
We propose an internal mechanism for LLMs to reduce extraneous load by re-organizing the positions of tokens. We formalize this process, termed context Re-Positioning (RePo), as learning non-constrained position values based on information relevance of tokens, instead of using the fixed linear positions in prior work. To this end, we introduce a differentiable module $f_{\phi}$, which assigns a position value in continuous space for each token based on its hidden state. The $f_{\phi}$ can be independently learned for each attention head of an LLM. Trained on general data, $f_{\phi}$ learns to re-position tokens free from conventional constraints like monotonicity or integer values. The continuity of modern position encoding functions, e.g., RoPE (Su et al., 2024) and ALiBi (Press et al., 2021) methods, is key to the end-to-end optimization of $f_{\phi}$, as it allows these assigned positions to be integrated in a differentiable manner.
@@
We find that LLMs trained with the RePo method achieve consistent performance gains on tasks involving noisy context, structured data, and long contexts. In our experiments, we continually pre-train LLMs using RePo and several baseline approaches based on the fully open-sourced OLMo-2 1B and 7B models, in order to avoid biased or misleading results arising from data contamination issues (Wu et al., 2025; Dong et al., 2024). On OLMo-2 1B, RePo outperforms RoPE by at least 5.4, 2.2, and 6.9 points on benchmarks targeting noisy context, structured data, and long context, respectively (Bai et al., 2024). We then evaluate RePo on OLMo-2 7B to assess its effectiveness at larger model scales. The results show similarly consistent performance gains, indicating that RePo scales well with increased model capacity. In addition to these improvements, RePo achieves comparable performance on a wide range of general benchmarks (Wang et al., 2024b; Clark et al., 2018; Zellers et al., 2019), which are typically short-context and require minimal reorganization.
@@
To better understand whether our method aligns with the CLT-based motivation and where the performance gains of the RePo method come from, we conducted a series of detailed analyses. First, to evaluate how different methods handle long-range dependencies, we compared the attention distributions across methods, particularly focusing on their treatment of distant but relevant information. In the NIAH task (Section 5.1), we observed that RePo allocates more attention to the most critical ``needle'' tokens compared to the baseline methods, while directing less attention to the nearest ``query'' tokens. This behavior breaks the typical locality bias (Yang et al., 2025b; Press et al., 2021), dynamically adjusting based on the context. Second, the positions assigned by RePo exist in a denser, more non-linear space, which is critical for enhancing its generalization to longer contexts, such as extending from 4K to 16K tokens. Finally, an interesting finding is that RePo learns positional patterns that resemble a hybrid of prior position-assignment strategies. These include assigning constant positions $a\pm\epsilon$, as in NoPE (Kazemnejad et al., 2023), and enforcing a monotonic position sequence, as in RoPE (Su et al., 2024) and other position encoding methods (Vaswani et al., 2017; Press et al., 2021), within a given context span (Section 5.2). In our case study (Appendix D), we also found that the positions assigned by RePo capture the intrinsic structure of the input context (e.g., segmentation of few-shot examples).
@@ 2 Background
The position information of tokens in a context is critical for the attention mechanism of LLMs. The position of a given token is generally mapped into embeddings (Su et al., 2024; Vaswani et al., 2017; Li et al., 2025) or biases (Press et al., 2021; Raffel et al., 2020) through a position encoding module before the attention function. In this section, we will introduce the notations for attention mechanism and position encoding functions.
@@
We use the rotary position encoding (RoPE) (Su et al., 2024) as a specific example to illustrate the position assignment and encoding procedure commonly employed in LLMs (Walsh et al., ; Yang et al., 2025a; Qwen et al., 2025). First, each token $x_{i}$ is assigned an integer position index $i$. This positional information is then incorporated into the model through an encoding function $g_{\theta}:\mathbb{R}\rightarrow\mathbb{R}^{d\times d}$, which is a differentiable function that generates a rotation matrix. The parameter $\theta$ represents a pre-defined frequency vector, generally frozen during training. In RoPE, the function $g_{\theta}$ is directly applied to the query $\bm{q}_{i}$ and key $\bm{k}_{j}$ as follows:
@@
In many related works, strict linear position assignment (i.e., 0 to L1L-1) similar to RoPE has become the standard approach, with various encoding functions being explored (Vaswani et al., 2017; Press et al., 2021; Yang et al., 2025b; Li et al., 2025). One notable exception is the NoPE method (Kazemnejad et al., 2023; Yang et al., 2025b), which omits both position assignment and position encoding. However, we demonstrate that this approach is equivalent to applying position encoding at a constant position $a$ for all tokens. Further discussion about the inter-connection between constant (e.g., NoPE) and linear position assignment (e.g., RoPE) strategies are provided in Appendix A.
@@ 3.1 Overview
The current linear context organization, which assigns consecutive integer indices to tokens, overlooks the internal structure of tokens based on their relevance. For instance, this limitation leads to noticeable performance degradation in tasks involving long-distance dependencies (Hsieh et al., ), a problem analogous to the high extraneous load issue in Cognitive Load Theory (CLT).
The main goal of this work is to reduce the unnecessary cognitive cost raised by the oversimplified organization of the context, i.e., extraneous cognitive load, thereby conserving the finite working memory capacity for beneficial germane load, such as the attention mechanism. To this end, we propose a context re-positioning (RePo) module $f_{\phi}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, which is a light-weight neural network that assigns more appropriate positions of tokens, taking into account their relevance within a given context. The assigned positions by $f_{\phi}$ are defined in a continuous, non-linear space, and can thus be optimized jointly with LLMs when equipped with differentiable position encoding methods (Vaswani et al., 2017; Su et al., 2024; Press et al., 2021). Notably, our RePo module $f_{\phi}$ is prior to position encoding, where the latter aims to map assigned positions into embeddings or biases.
@@ 3.2 Context Re-positioning
Kazemnejad et al. (2023) shows that position information may be entangled in original hidden states, so the first component is designed to extract the position representation from hidden states of tokens explicitly. In our implementation, we use a light-weight SwiGLU (Shazeer, 2020) sub-layer to achieve this goal:
@@
The subsequent component assigns a new position value $z_{i}$ for token $x_{i}$ in each attention head. There are a variety of modeling strategies for processing $\bm{r}_{i}$ with full (Vaswani et al., 2017) or limited (LeCun and Bengio, 1998) access to $\bm{r}_{<i}$, but we find that a linear transformation achieves comparable performance with lower latency:
@@
where the position encoding function $g_{\theta}$ is the same as that used in RoPE in Eq. 3. It is worth noting that our RePo is not restricted to RoPE and can be easily extended to all the differentiable position encoding methods (Vaswani et al., 2017; Press et al., 2021; Li et al., 2025). In practice, we apply the position representation module (Eq. 4) for each layer independently, and position assignment module (Eq. 5) for each attention head independently.
Table 1: Performance on noisy context. We evaluate on needle-in-a-Haystack (NIAH) tasks in RULER benchmark (Hsieh et al., ) with four types of needles: single (S), multi-query (MQ), multi-key (MK), and multi-value (MV). The best results are in bold text.
Base Method S MQ MK MV AVG. $\Delta$ vs RoPE
1B RoPE 100.0 78.5 86.0 79.0 85.9 0.0
NoPE 83.0 83.3 76.0 78.5 80.2 -5.7
R2N1 100.0 91.8 82.0 88.3 90.5 ++4.6
N2R1 95.0 89.8 84.0 84.3 88.3 ++2.4
RePo 100.0 89.3 86.0 89.8 91.3 ++5.4
7B RoPE 100.0 94.8 97.0 97.3 97.3 0.0
RePo 100.0 95.5 96.0 100.0 97.9 ++0.6
@@
Since many position encoding methods are differentiable (Su et al., 2024; Press et al., 2021; Li et al., 2025), we can train an LLM with RePo-based attention (Eq. 7) using backpropagation when equipped with such encodings. To balance efficiency and effectiveness, we apply the RePo method starting from the $1/3$-th layer of LLMs while keeping standard position encoding for the lower layers. This design choice is motivated by previous findings that the lower layers of LLMs primarily capture surface-level features that depend more on local information (Tenney et al., 2019), such as part-of-speech tagging and syntax, and thus benefit less from reorganization. An ablation study for the hyper-parameter is in Appendix B.5.
@@ 4 Experiments
Figure 1: Long-context evaluation on OLMo-2 1B model. We use the data in RULER (Hsieh et al., ) benchmark and apply YaRN (Peng et al., 2024) for all RoPE layers to extend the context. Similar results on OLMo-7B are shown in Appendix B.4.
@@ 4.1 Setup
We use OLMo-2 (Walsh et al., ) as the backbone model, which is fully open-sourced, including its training data, model weights, and code. Our primary motivation for conducting experiments on a fully open-source LLM is to avoid data contamination issues (Wu et al., 2025; Dong et al., 2024), which can otherwise lead to biased or misleading results. Notably, OLMo-2 is trained with the RoPE method with linear position assignment (Su et al., 2024). For all methods in our experiments, we start from the checkpoint of the OLMo-2 1B and 7B model that have completed stage-1 pre-training on 4 trillion tokens, and we continually pre-train it on the 5050B-token stage-2 data [^3] https://github.com/allenai/OLMo. The training context length is 4096 tokens. We keep the training configuration and codebase identical to those released by Walsh et al. .
Table 2: Performance for structured data. We evaluate on HybridQA dataset (Chen et al., 2020) for table data. We use exact match (EM) as metric. We highlight the best results with bold text.
Base Method Table $\Delta$ vs RoPE
1B RoPE 24.43 0.00
NoPE 23.52 -0.91
R2N1 25.11 ++0.68
N2R1 23.86 -0.57
RePo 26.70 ++2.27
7B RoPE 33.52 0.00
RePo 37.61 ++4.09
@@
For our RePo method, we use the RoPE function (Su et al., 2024) to encode the dynamic positions assigned by $f_{\phi}$, as shown in Eq. 7. We apply it starting from the $1/3$-th layer of the OLMo-2 model, i.e., the 5th and 10th layers for the 1B and 7B models, respectively. In each layer that uses RePo, we share the parameters for the position representation transformation in Eq. 4, while learning the position assignment for each attention head independently, as described in Eq. 5. The hidden size of the learned position representation is set to $1/8$ of the model’s hidden size, based on the assumption that positional information is less rich than the information in the original hidden states. As shown in §5.4, RePo is lightweight and introduces negligible overhead compared to the original LLM.
We compare RePo with followings on OLMo-2 1B:
  • RoPE: Uses RoPE (Su et al., 2024) for positional encoding, identical to pre-training the process (Walsh et al., ). Unless otherwise specified, this method uses linear position assignment during training, i.e., from 0 to 4095.
@@
  • NoPE: Removes explicit positional encoding methods (Kazemnejad et al., 2023; Wang et al., 2024a), i.e., RoPE is omitted during continual pre-training. This is equivalent to constant position assignment (Appendix A).
@@
  • R2N1: A hybrid positional encoding method that interleaves RoPE and NoPE (Yang et al., 2025b; Meta, 2025). For every three layers, the first two use RoPE, while the last one uses NoPE.
  • N2R1: The opposite baseline of R2N1.
@@
We further evaluate RePo on the OLMo-2 7B model using RoPE as a baseline to demonstrate the effects of scaling up the model size. We train those models on 4 H100 GPUs for 50B tokens. We use the allenai/olmes [^4] https://github.com/allenai/olmes codebase for evaluation, which provides extensive test suites. We categorize our main evaluation tasks into three dimensions:
@@
  • Noisy Context evaluates the model’s robustness to contexts containing large amounts of irrelevant information. Such ``noise’’ increases extraneous cognitive load (Paas et al., 2003), which can negatively affect problem-solving performance (Peysakhovich and Lerer, 2023). We use multiple variants of needle-in-a-hystack (NIAH) tasks provided by RULER benchmark (Hsieh et al., ) for evaluation, which purposefully constructs contexts with irrelevant content for evaluation.
@@
  • Structured Data evaluates performance on structured data, e.g., tables. This setting highlights the importance of contextual organization, as linearizing such data into natural language often leads to significant structural information loss. We use HybridQA (Chen et al., 2020) to evaluate the performance on table reasoning.
@@
  • Longer Context evaluates model performance on sequences longer than those seen during training (i.e., 4K tokens), as positional encoding has been shown to strongly affect long-context generalization (Press et al., 2021). We use subsets of RULER and LongBench (Bai et al., 2024), which contain examples with 4K to 16K tokens.
@@
Notably, both the evaluations for noisy context and structured data are conducted within the training context length (i.e., 4K tokens). In contrast, the longer-context evaluation requires extended context lengths. To enable this, we apply the YaRN method to RoPE layers across all methods using the extrapolation hyperparameters in Peng et al. (2024). More details are in Appendix B.
@@
Table 3: Performance on LongBench. This uses F1 score for QA and fewshot tasks, while Rouge-L (Lin, 2004) for the summarization tasks. We only evaluate on data that contains less than 16K tokens. Since the model is trained with a max of 4096 tokens, the YaRN method (Peng et al., 2024) is used for all RoPE layers for context extrapolation. We use bold for best results.
@@
Base Method Multidoc QA Singledoc QA Summarization Fewshot AVG. $\Delta$ vs RoPE
2WikiMultihopQA MuSiQue MultiFieldQA-en NarrativeQA GovReport QMSum TriviaQA
1B RoPE 23.32 7.24 27.37 12.94 6.23 7.96 61.47 20.93 0.00
NoPE 9.11 0.33 13.64 1.80 5.15 0.68 18.12 6.98 -13.95
R2N1 25.88 7.31 31.28 16.24 4.53 8.74 66.67 22.95 ++2.02
N2R1 16.24 0.40 21.88 1.26 8.74 5.31 25.18 11.29 -9.64
RePo 30.86 13.45 33.12 15.24 16.80 12.53 73.02 27.86 ++6.93
7B RoPE 22.10 6.69 32.61 16.38 11.19 6.63 83.97 25.65 0.00
RePo 33.41 19.55 40.37 22.64 12.30 10.41 85.50 32.03 ++6.38